# On closure operations in the space of subgroups and applications

### v1

@@ 5.1. Proofs of the results
Note that if a set of laws detects amenability in a group $G$, it also detects amenability in any subgroup of $G$. The following is an immediate consequence of the definition and Lemma 2.9.
@@
When the group $G$ is residually finite, there is a shorter way to obtain the conclusion of Theorem 5.12. Indeed, since the $G$-space $\partial_{F}G$ is proximal, it is hereditarily minimal [Gla76]. Moreover it follows from the conclusion of Proposition 2.1 that being hereditarily minimal is inherited from a $G$-space to its stabilizer URS. Hence $\mathcal{A}_{G}$ is a hereditarily minimal URS. Hence Propositions 4.11 and 4.12 apply, and the conclusion follows as above.

### v3

@@ 5.1. Proofs of Proposition 4 and Theorem 1
We deduce Proposition 4 from the introduction, that we state again for the reader’s convenience:
@@
Note that if a set of laws detects amenability in a group $G$, it also detects amenability in any subgroup of $G$.
Recall that a group $G$ satisfies the Tits alternative if every subgroup of $G$ is either virtually solvable or contains a non-abelian free subgroup. Beyond the original result of Tits about linear groups [Tit72], many classes of groups are known to satisfy the Tits alternative. Examples include hyperbolic groups [Gro87], the outer-automorphism group of a finitely generated free group [BFH00, BFH05] and of many free products [Hor14], groups acting properly on a finite-dimensional CAT(0) cube complex and with finite subgroups of bounded order [SW05], the group of polynomial automorphisms of $\mathbb{C}^{2}$ [Lam01], the Cremona group of birational transformations of the projective space $\mathbb{P}^{2}(\mathbb{C})$ [Can11, Ure21].
Proposition 5.11.
If a group $G$ satisfies the Tits alternative, then there is a set of laws that detects amenability in $G$.
Proof.
For $\ell\geq 1$, let $w_{\ell}$ be the law from Example 2.9. So a group is solvable of length at most $\ell$ if and only if it satisfies $w_{\ell}$. Let $H$ be a subgroup of $G$. Since $G$ satisfies the Tits alternative, $H$ is amenable if and only if $H$ is virtually solvable, if and only if there is $\ell$ such that $H$ virtually satisfies $w_{\ell}$. So $\mathbb{W}=\left\{w_{\ell}\right\}_{\ell\geq 1}$ detects amenability in $G$. ∎
The following is an immediate consequence of the definition and Lemma 2.10.
@@
We note that in view of Proposition 5.11, the following result covers Theorem 1 from the introduction. And this result applies to all the groups (and to all their subgroups) satisfying the Tits alternative mentioned above.
@@
When the group $G$ is residually finite, there is a shorter way to obtain the conclusion of Theorem 5.13. Indeed, since the $G$-space $\partial_{F}G$ is proximal, it is hereditarily minimal [Gla76, Lemma 3.2]. Moreover it follows from the conclusion of Proposition 2.1 that being hereditarily minimal is inherited from a $G$-space to its stabilizer URS. Hence $\mathcal{A}_{G}$ is a hereditarily minimal URS. Hence Proposition 4.11 and Theorem 4.12 apply, and the conclusion follows as above.